A swimming pool is in the shape of a rectangle. The length of the pool is 20 feet more than the width. A concrete deck 6 feet wide is added around the pool, and the perimeter of the deck is 160 feet. Find the dimensions of the swimming pool.
Answer by ankor@dixie-net.com(22740)

A swimming pool is in the shape of a rectangle.
The length of the pool is 20 feet more than the width.
A concrete deck 6 feet wide is added around the pool, and the perimeter of the deck is 160 feet.
Find the dimensions of the swimming pool.
:
L = W + 20
:
A 6' wide deck around the pool adds 12' to each dimension
2(L+12) + 2(W+12) = 160
Simplify, divide equation by 2
L + 12 + W + 12 = 80
L + W + 24 = 80
L + W = 80 - 24
L + W = 56
Replace L with (W+20)
W+20 + W = 56
2W = 56 - 20
2W = 36
W = 36/2
W = 18' is the width of the pool
therefore
18 + 20 = 38' is length
:
:
See if that checks out, find the outside perimeter with these values
2(38+12) + 2(18+12) =
2(50) + 2(30) =
100 + 60 = 160